OpenStudy (anonymous):

A rabbit and a turtle were to start a race at the same place. The rabbit was asleep when the turtle started. He slept for seven hours before he started the race. The rabbit's average rate when he was awake was 5 km/hr. The turtle never stopped for a break, but his pace was only 1km/hr. How long will the rabbit run before it catches the turtle?

OpenStudy (anonymous):

r* t = d Let t in hours be the time that the turtle runs to catch the rabbit. The turtle is 5hr * 1 km/hr or 5km in front of the rabbit. Let a stop watch start when the rabbit bolts for the turtle. Solve the following equation for t:\[5\text{km}+1 \frac{\text{km}}{\text{hr}}*(t)=t\left(5\frac{\text{km}}{\text{hr}}\right)\]\[t=\frac{5}{4} \text{hr} \text{ or } 1\text{ hr } 25 \text{ minutes} \]

OpenStudy (anonymous):

Sorry 5/4 hr = 1 hr 15 minutes